# Multi-Level Explanations for Generative Language Models

### v1

@@ 3.1 Handling Text Outputs
[…]
  • Log NLI: Uses a natural language inference (NLI) model to predict the log-odds of entailment given $y^{o}$ as premise and $y$ as hypothesis, and optionally in the other direction as well (see Appendix A.1).
3.2 Handling Long Text Inputs
Some generative language tasks require long input texts, such as in summarization and context-grounded QA. We address this challenge through a combination of three techniques: segmenting the input into linguistic units at multiple levels, using attribution algorithms with linear complexity in the number of input units, and obtaining attributions to units in a coarse-to-fine manner.
Linguistic segmentation
We segment the input text into linguistic units at multiple granularities: paragraphs, sentences, phrases, and words (“Select Units” box in Figure 2). This approach takes advantage of linguistic and other structure present in the input. For example, the input may already be broken into paragraphs or contain multiple distinct retrieved documents, in which case these paragraphs or documents can form the units at the highest level. In contrast, many existing methods rely on the model’s tokenizer, which can yield units (tokens) that are too fine, or treat the text as a flat sequence of tokens and let the algorithm decide how to segment it [2, 25].
We use tokenization and dependency parsing from spaCy v3.6 [9] to segment paragraphs into sentences and words. To segment sentences into phrases, we propose an algorithm that uses the dependency parse tree from spaCy. In the first pass, the algorithm recursively segments the tree and its subtrees into phrases that are no longer than a maximum phrase length. In the second pass, some short phrases are re-merged. More details are in Appendix A.2.
Our framework allows for units at any level to be marked as not of interest for attribution. For example at the word or token level, punctuation and stop words may be ignored. In question answering, prompts for LLMs often follow a template, e.g. prefacing the context with the string “Context: ”. Such elements of the template are usually not of interest. We may also not attribute to the question and focus only on the context.
Linear-complexity algorithms
Given an input segmentation (which may have mixed units as in Figure 2), we have the actual task of attributing to units $x_{1},\dots,x_{d}$ (the “Post Hoc Explainer” block in Figure 2). For this we consider only attribution algorithms that scale linearly with the number of units $d$ in terms of model queries, to control this cost. We instantiate MExGen with three such algorithms: Local Shapley (L-SHAP), a LIME-like algorithm with further constraints (C-LIME), and leave-one-out (LOO) as a baseline.
L-SHAP: This local variation of SHAP was proposed by Chen et al. [4] to decrease the number of model inferences relative to SHAP, which requires exponentially many inferences. In our context, L-SHAP does so by only perturbing units that are within a constant-size neighborhood of the current unit being attributed to. This makes the number of inferences scale linearly with the number of units. More precisely, for unit of interest $s\in[d]$, we consider only the radius-$M$ neighborhood $s-M,\dots,s+M$ (truncated to $1,\dots,d$ if necessary), and define $\Omega^{M,K}_{s}$ to be the set of all subsets of $s-M,\dots,s+M$ up to cardinality $K$. Then the attribution score $\xi_{s}$ for unit $s$ is given by
$\xi_{s}$ $=\sum_{A\in\Omega^{M,K}_{s}}\frac{S(x^{(A)};y^{o},f)-S(x^{(A\cup\{s\})};y^{o},f)}{Z_{|A|}},$ (2)
where $x^{(A)}$ is a perturbation of $x^{o}$ in which units $j\in A$ are replaced (e.g., by a baseline value such as [MASK] or using a LM), and $Z_{|A|}$ is the number of subsets in $\Omega^{M,K}_{i}$ with cardinality $|A|$.
C-LIME: Similar to LIME, we use a linear model that operates on interpretable features $z$ and approximates the model $f$ in the vicinity of original input $x^{o}$. In our case, the interpretable features $z\in\mathbb{R}^{d}$ correspond to the units $x_{1},\dots,x_{d}$, with $z_{s}=0$ if unit $s$ is perturbed and $z_{s}=1$ otherwise. The linear model is fit using perturbations $x^{(1)},\dots,x^{(n)}$ of $x^{o}$, with corresponding interpretable representations $z^{(1)},\dots,z^{(n)}$, and model outputs (scalarized in our case) $S(x^{(1)};y^{o},f),\dots,S(x^{(n)};y^{o},f)$:
$\xi=\operatorname*{arg\,min}_{w}\sum_{i=1}^{n}\pi\bigl{(}z^{(i)}\bigr{)}\bigl{(}w^{T}z^{(i)}-S(x^{(i)};y^{o},f)\bigl{)}^{2}+\lambda R(w),$ (3)
We make two main departures from LIME, in addition to the use of a general scalarizer $S$ as seen in (3). Firstly, we limit the number of perturbations $n$ to a multiple of the number of units $d$. Since the regression parameters $w$ are $d$-dimensional (not including a possible intercept), the multiplier could be set to $10$ for example to have $10$ times as many samples (i.e., perturbations) as parameters to fit. In contrast, LIME by default sets $n$ to be in the thousands independently of $d$, which can be prohibitive for LLMs. Secondly, we limit the number of units $K$ that can be perturbed simultaneously to a small integer. This concentrates the smaller number of perturbations on inputs that are closer to $x^{o}$, which has been shown to improve the local fidelity of attributions [30]. LIME by contrast samples the number of units to perturb uniformly from $\{1,\dots,d\}$ [15]. Other details are in Appendix A.3.
LOO: We also consider a LOO variant of MExGen as a baseline. Here, units $x_{1},\dots,x_{d}$ are perturbed one at a time to yield perturbations $x^{(1)},\dots,x^{(d)}$ (i.e., $n=d$). The attribution score for $x_{s}$ is the corresponding decrease in scalarizer score from its original value, $\xi_{s}=S(x^{o};y^{o},f)-S(x^{(s)};y^{o},f)$.
Multi-level explanations
Once we have computed attribution scores at a given level, we may choose to refine the input units and repeat the process (feedback path at the bottom of Figure 2). For example, given sentence-level attributions for a long input, we may wish to obtain phrase- or word-level attributions for a few sentences and keep attributions at the sentence level for the remainder to not introduce too many new units. The few sentences can be selected by any number of automatic criteria or by the user, or the user can even choose to stop at sentence level. See Appendix B for sentence selection criteria used in our experiments.
@@ 4 Automated Evaluation
We evaluate MExGen across two text generation tasks: Context-grounded QA and summarization.
@@ 4.1 Setup
Datasets
For summarization, we use two popular datasets, Extreme Summarization (XSUM) [19] and CNN/Daily Mail (CNN/DM) [23, 8]. For question answering, we use the Stanford Question Answering Dataset (SQuAD) [21] because of its relatively short, paragraph-long contexts, which allow us to test mixed sentence-and-word-level attributions.
Models
For summarization, we use two models, a 306M-parameter distilbart-xsum-12-6 model [^3] https://huggingface.co/sshleifer/distilbart-xsum-12-6, and a larger 20B-parameter Flan-UL2 model [31]. We treat the DistilBart model as one with full logit access, thus enabling use of the Log Prob scalarizer. We call Flan-UL2 through an API. [^4] We use Flan-UL2 via the IBM Generative AI Python SDK (Tech Preview) [10]. For QA, we use the 770M-parameter Flan-T5-Large model [^5] https://huggingface.co/google/flan-t5-large [5] and treat it as providing output logits.
@@
We instantiate MExGen with the attribution algorithms discussed in Section 3.2: L-SHAP, C-LIME, and LOO, where the third is intended more as a baseline. For summarization, we obtain mixed sentence and phrase attributions while for QA, we obtain mixed sentence and word attributions. Algorithm parameter settings are given in Appendix B. We compare against PartitionSHAP (P-SHAP) [25] and CaptumLIME [16]. For PartitionSHAP, recall from Section 2 that it requires the selection of an output token to explain, whereas we are interested in the output sequence as a whole. For this reason, we slightly modify PartitionSHAP by summing across all attribution scores, corresponding to different output tokens, that it gives to each input span. This is equivalent to explaining the sum of the log probabilities of output tokens because of the linearity of Shapley values, and it corresponds to our Log Prob scalarizer. For CaptumLIME, since it accepts user-defined units for attribution, we provide it with the same sentence-level and mixed-level units used by our algorithms. Thus we can directly compare Captum’s attribution algorithm (i.e., LIME) with ours, controlling for input segmentation. Captum’s default for the target function to explain also corresponds to the Log Prob scalarizer. Additionally, PartitionSHAP and CaptumLIME take the number of model queries as an input; for a fair comparison, we allow them the greater of the numbers of model calls used by MExGen L-SHAP and C-LIME.
@@
Figure 3: Spearman’s rank correlation between attribution scores using different scalarizers. Attributions were computed using multi-level (a) MExGen C-LIME for the DistilBart model on the XSUM dataset and (b) MExGen L-SHAP for Flan-T5-Large on the SQuAD dataset.
Metrics
To measure local fidelity of explanations, we use perturbation curves as in previous work on explaining text classification [2, 12]. The general idea behind perturbation curves is to rank input units in decreasing importance according to a set of attribution scores, perturb the top $k$ units with $k$ increasing, and plot the resulting change in some output score that measures how much the perturbed output deviates from the original output. For the output score, we select an evaluation scalarizer from the set of scalarizers in Section 3.1. To accommodate PartitionSHAP and CaptumLIME, we choose an evaluation scalarizer corresponding to the target function that they use for explanation. For the DistilBart summarization and Flan-T5-Large QA models, this means the Log Prob scalarizer, while for Flan-UL2, we choose SUMM in keeping with PartitionSHAP. Appendix C has further details on how we compute perturbation curves, accounting for the different units used by different attribution algorithms.
4.2 Scalarizer Evaluation
Unit Ranking Similarity Across Scalarizers
We evaluate the similarity between attribution score vectors $\xi_{S}$ and $\xi_{S^{\prime}}$ at the same granularity level for all pairs of scalarizers $S,S^{\prime}$. We present Spearman rank correlation to measure similarity in unit rankings. See Appendix C for further results.
Figure 3 shows Spearman correlation matrices, averaged across examples from the respective datasets. Certain pairs are highly similar, for example BART and SUMM as mentioned in Section 3.1, as well as Sim and BERT. Log Prob, the only one that uses logits from the model being explained, clearly differs from the others. There are also noticeable differences between panels (a) and (b) in Figure 3, which feature different tasks, datasets, and models. For example, Spearman correlation between BERT and SUMM is $0.75$ in (a) and $0.83$ in (b). This suggests the necessity of exploring different scalarizers to explain different tasks.
(a) BERT Score as Evaluation
(b) Log Prob as Evaluation
(c) BART Score as Evaluation
Figure 4: Perturbation curves for MExGen C-LIME with different scalarizers, used to explain the DistilBart model on the XSUM dataset. The curves show the decrease in (a) BERTScore, (b) log probability, and (c) BARTScore when removing the most important $p$of tokens according to each explanation scalarizer. Shading shows standard error.
@@ 4.2 Scalarizer Evaluation
Figure 4 shows perturbation curves with different evaluation scalarizers for MExGen C-LIME using different scalarizers to compute attribution scores (i.e., a cross-evaluation of scalarizers). Results are again averaged over examples from the datasets (see Appendix C for how). In panels (b) and (c), the case where the attribution scalarizer is matched with the evaluation scalarizer is generally the best. Additionally, the Log Prob scalarizer performs well across all three evaluation scalarizers. This implies that the Log Prob scalarizer gives the most universal (least dependent on the evaluation scalarizer) explanation; hence, when log probabilities are available, we suggest using them.
In the case where model logits are not available (e.g., when only API access is given), the choice of scalarizer is not clear. However, if (hypothetically) one were to use Log Prob as the evaluation scalarizer, our results in Figures 3 and 4(b) indicate that BERT score is the scalarization method that best approximates explanations with Log Prob as the scalarizer. We include results for other tasks, models, and evaluation scalarizers in Appendix C.
@@ 4.3 Comparison Between Explainers
(a) Flan-UL2 on CNN/DM
(b) DistilBart on XSUM
@@
Figure 5: Perturbation curves for different explanation methods, using the same scalarizer as the evaluation indicated on the y-axis. Models and datasets: (a) Flan-UL2 on the CNN/DM dataset, (b) DistilBart on XSUM, and (c) Flan-T5-Large on SQuAD. MExGen provides multi-level explanations down to phrase-level for XSUM and CNN/DM and word-level for SQuAD. We also show the perturbation curve of MExGen C-LIME using mismatched scalarizers: BARTScore (a) and BERTScore (b)(c). Shading shows standard error in the mean.
@@
We compare the performance of the MExGen instantiations with P-SHAP and CaptumLIME. We were only able to generate results for CaptumLIME in Figure 5 (b) because (i) we had API access to Flan-UL2 and CaptumLIME needs output logits, and (ii) CaptumLIME does not support Flan-T5-Large. Figure 5 shows that perturbing units (i.e., the tokens that they comprise) ranked as most important by MExGen has the largest influence on the evaluation scalarization, compared to P-SHAP and CaptumLIME. This indicates that the methods using MExGen identified more important tokens according to the evaluation scalarizer, while P-SHAP and CaptumLIME give higher scores for less important tokens.
We found that MExGen C-LIME consistently outperforms the rest of the methods across all tested models and datasets. In particular, the curve for MExGen C-LIME is always higher than that of P-SHAP up to at least the top 10% of tokens; the gray vertical lines in Figure 5 show the fraction of tokens where the cross-over occurs. We also show in Figure 5 the perturbation curve for MExGen C-LIME using a different scalarizer than used for evaluation. Surprisingly, we find that even when using a mismatched scalarizer (notably BERT in (b)(c), which does not even use logits), MExGen C-LIME can outperform P-SHAP in local fidelity.
@@
XSUM DistilBart
14.7 14.7 14.1 10.7
Flan-UL2 21.3 21.2 18.4 20.0
CNN/DM DistilBart
13.5 14.7 13.2 9.7
Flan-UL2 31.7 30.8 28.3 30.7
SQUaD Flan-T5-Large 61.2 59.9 59.9 59.4
Table 2: Area under the perturbation curve (AUPC, higher is better) up to $20$% of tokens. For DistilBart and Flan-T5-Large, log probability is used as both the explanation and evaluation scalarizer. For Flan-UL2, which does not provide access to log probabilites, SUMM is used as the explanation and evaluation scalarizer.
@@
Table 2 shows the area under the perturbation curve (AUPC) to summarize performance over all dataset-model pairs that we tested, not just those in Figure 5. We evaluate AUPC up to $p=20\%$ as done in Chen et al. [2]. Across all dataset-model pairs, methods that use the MExGen framework (sometimes including the baseline LOO) performed better (i.e., higher AUPC) than P-SHAP. Moreover, the second highest AUPC is always a MExGen method.

### v2

@@ 3.2 Handling Long Text Inputs
6:for  $s\in[d]$ do
7:     if $\psi_{s}\geq\phi\textbf{ and }\psi_{s}\in\text{Top-}k(\psi)$ then
8:         units_to_be_refined.add($x_{s}$)
9:return units_to_be_refined
@@ 4 Automated Evaluation
We evaluate MExGen on two text generation tasks, summarization and context-grounded QA. Section 4.2 compares the scalarizers that we have considered and provides guidance on the choice of scalarizer. Sections 4.3 and 4.4 compare MExGen to other attribution algorithms and to LLM self-explanations respectively, showing that MExGen has higher fidelity in terms of identifying input units that are most important to the explained model.
@@ 4.1 Setup
Datasets and LMs
For summarization, we evaluate on the Extreme Summarization (XSUM) Narayan et al. (2018) and CNN/Daily Mail (CNN/DM) See et al. (2017); Hermann et al. (2015) datasets. In Sections 4.2 and 4.3, we use three LMs: the 306M-parameter DistilBART  [^4] https://huggingface.co/sshleifer/distilbart-xsum-12-6, the 20B-parameter Flan-UL2 Tay et al. (2023), and the Llama-3-8B-Instruct (“Llama-3”) (Dubey et al., 2024) models. We treat the DistilBART model as one with full logit access, enabling use of the Log Prob scalarizer. We call Flan-UL2 and Llama-3 through an IBM API (IBM, 2024) that only returns text, thus representing the text-only setting precluding the use of Log Prob. We evaluate on the first 1000 test set examples of each dataset for DistilBART and the first 500 for Flan-UL2 and Llama-3.
For QA, we use the Stanford Question Answering Dataset (SQuAD) Rajpurkar et al. (2016) (1000 validation set examples, see Appendix C.1). We consider two LMs, Llama-3 through the above API, and the 770M-parameter Flan-T5-Large model [^5] https://huggingface.co/google/flan-t5-large Chung et al. (2022), which we treat as providing output logits.
@@
We instantiate MExGen with the attribution algorithms discussed in Section 3.2: LOO, C-LIME, and L-SHAP. For summarization, we obtain mixed sentence- and phrase-level attributions while for QA, we obtain mixed sentence and word attributions, which are appropriate for SQuAD’s shorter, paragraph-long contexts. Choices for scalarizers, segmentation, and algorithm parameters are described in Appendix C.3.
We compare against PartitionSHAP (P-SHAP) SHAP (2024a) and CaptumLIME Miglani et al. (2023). For PartitionSHAP, recall from Section 2 that it requires the selection of an output token to explain, whereas we are interested in the output sequence as a whole. For this reason, we modify PartitionSHAP by summing across all attribution scores, corresponding to different output tokens, that it gives to each input span. This is equivalent to explaining the sum of the log probabilities of output tokens because of the linearity of Shapley values, and it corresponds to our Log Prob scalarizer. For CaptumLIME, since it accepts user-defined units for attribution, we provide it with the same units used by MExGen. Thus we can directly compare Captum’s attribution algorithm (i.e., LIME) with ours, controlling for input segmentation. Captum’s default for the target function to explain also corresponds to the Log Prob scalarizer. Additionally, PartitionSHAP and CaptumLIME take the number of model queries as an input; for a fair comparison, we allow them the greater of the numbers of model calls used by MExGen L-SHAP and C-LIME.
Metrics
To measure local fidelity of explanations, we use perturbation curves as in Chen et al. (2020); Ju et al. (2023). Given a set of attribution scores, perturbation curves rank input units in decreasing importance according to their scores, perturb the top $k$ units with $k$ increasing, and plot the resulting change in some output scalarization measuring how much the perturbed output deviates from the target output. For the output scalarization, we select an evaluation scalarizer from those in Section 3.1. To accommodate PartitionSHAP and CaptumLIME (and slightly favor them), we choose an evaluation scalarizer corresponding to the target function that they use for explanation. For DistilBART and Flan-T5-Large, this means the Log Prob scalarizer, while for Flan-UL2 and Llama-3, we choose SUMM in keeping with PartitionSHAP. Please see Appendix C.5 for further details.
4.2 Scalarizer Evaluation
Ranking Similarity Across Scalarizers
We first compare scalarizers by measuring the Spearman rank correlation between attribution score vectors $\xi_{S}$ and $\xi_{S^{\prime}}$ at the same granularity for all scalarizer pairs $S,S^{\prime}$. Results with cosine similarity instead of Spearman are in Appendix D.1, Figure 7.
Figure 3 shows Spearman correlation matrices, averaged across examples from the respective dataset. Certain scalarizer pairs are highly similar, for example BART and SUMM as mentioned in Section 3.1, as well as Sim and BERT. Thus, it may suffice to consider only one from each pair. Log Prob, the only one that uses logits from the model being explained, clearly differs from the others. Additional results in Figure 7 show the same patterns. There are also noticeable differences between Figures 3a and 3b, which feature different tasks, datasets, and models. For example, Spearman correlation between BERT and SUMM is $0.75$ in (a) and $0.83$ in (b). This suggests the necessity of exploring different scalarizers for different tasks.
@@ 4.2 Scalarizer Evaluation
Figure 3: Spearman’s rank correlation between attribution scores using different scalarizers. Attributions were computed using multi-level (a) MExGen C-LIME for the DistilBART model on the XSUM dataset and (b) MExGen L-SHAP for Flan-T5-Large on SQuAD.
(a) Log Prob as Evaluation
(b) BERT Score as Evaluation
(c) SUMM Score as Evaluation
Figure 4: Perturbation curves (higher is better) for MExGen C-LIME with different scalarizers, used to explain the DistilBART model on the XSUM dataset. The curves show the decrease in (a) log probability, (b) BERTScore, (c) SUMM score when removing the most important $p$of tokens according to each explanation scalarizer. Shading shows standard error in the mean.
@@
Figures 4a4c show perturbation curves with different evaluation scalarizers, using MExGen C-LIME with all proposed scalarizers (i.e., a cross-evaluation of scalarizers). The curves are averaged over 1000 samples from XSUM (see Appendix C.5 for how) and the shading shows one standard error above and below. The case where the attribution scalarizer is matched with the evaluation scalarizer is generally the best. The Log Prob scalarizer performs well across evaluation scalarizers, implying that it gives a more universal explanation (least dependent on the evaluation scalarizer). Hence, we suggest using the log probabilities when they are available.
When model logits are not available, the choice of scalarizer is not as clear. However, if (hypothetically) one were to use Log Prob as the evaluation scalarizer, Figures 3 and 4a indicate that the BERT scalarizer best approximates the Log Prob scalarizer (Sim is similar while SUMM, the scalarizer used by P-SHAP, performs worse in this regard). Results for MExGen L-SHAP and a second model-dataset pair in Appendix D.1, Figures 810 show similar trends as above.
The computational cost of different scalarizers is discussed in Appendix C.3.
@@ 4.3 Comparison Between Explainers
(a) DistilBART on XSUM
(b) Llama-3 on CNN/DM
@@
Figure 5: Perturbation curves (higher is better) for different explanation methods, using the same scalarizer as the evaluation scalarizer in the y-axis label. Models and datasets: (a) DistilBART on XSUM, (b) Llama-3 on CNN/DM, (c) Flan-T5-Large on SQuAD. The legend in (c) applies to all panels, but we could run CaptumLIME only for (a). “MExGen C-LIME (mismatched)” refers to the use of mismatched scalarizers: BERTScore (a)(c) and BARTScore (b).
@@
Figure 5 compares the perturbation curves of MExGen instantiations with P-SHAP and CaptumLIME. Mean curves and standard errors are again computed over the number of examples taken from each dataset (see Section 4.1).
Regarding CaptumLIME, we were only able to obtain results for it in Figure 5a because (i) we had API access to Llama-3 and CaptumLIME needs output logits, and (ii) CaptumLIME does not support Flan-T5-Large. Figure 5a is a direct comparison between LIME (represented by CaptumLIME) and our modification C-LIME, using the same number of model queries and input segmentation as mentioned earlier. C-LIME is clearly more effective.
In comparison with P-SHAP, Figure 5 shows that the perturbation curves for MExGen rise more quickly from zero and are higher for the top $x$of tokens, where $x>20$% in Figures 5a and 5b, and $x$ varies between 8% and 13% in Figure 5c. This pattern indicates that MExGen is better able to identify units that are most important to the model, as measured by the change in the evaluation scalarizer in the leftmost region of each plot. Figure 11 has results on additional model-dataset pairs. Across the tested models and datasets, MExGen C-LIME and L-SHAP are the top performers, while MExGen LOO, the simplest attribution algorithm, is close behind.
We also show in Figure 5 perturbation curves for MExGen C-LIME using different scalarizers than used for evaluation. Surprisingly, we find that even when using a mismatched scalarizer (notably BERT in Figures 5a, 5c, which does not even use logits), MExGen C-LIME can outperform P-SHAP in fidelity.
@@
XSUM DistilBART 13.6 13.8 13.1 9.4
Flan-UL2 17.2 17.4 16.7 13.7
Llama-3-8B-Instruct 22.4 22.2 22.1 20.2
CNN/DM DistilBART 13.5 14.7 13.2 9.7
Flan-UL2 32.1 32.0 32.1 33.2
Llama-3-8B-Instruct 26.4 26.3 26.1 22.1
SQuAD Flan-T5-Large 62.7 61.1 60.2 58.8
Llama-3-8B-Instruct 56.4 57.0 54.9 38.5
Table 1: Areas under the perturbation curve (AUPC) up to $20$% of tokens. For DistilBART and Flan-T5-Large, log probability is used as both the explanation and evaluation scalarizer. For Flan-UL2 and Llama-3-8B-Instruct, which do not provide access to log probabilites, SUMM is used as the explanation and evaluation scalarizer.
@@
Table 1 shows the area under the perturbation curve (AUPC) to summarize performance over all dataset-model pairs that we tested, including the ones in Figure 11. We evaluate AUPC up to $20\%$ of tokens as done in Chen et al. (2020). Across all dataset-model pairs, MExGen instantiations (including LOO) performed better (i.e., higher AUPC) than P-SHAP. The one exception is the (CNN/DM, Flan-UL2) pair, but even in this case, Figure 11b in Appendix D.2 shows that the MExGen curves are higher for the top 5% of tokens. The second highest AUPC is always from MExGen C-LIME or L-SHAP.
Computational Cost
The computational cost of the compared explanation methods — measured by time and memory use — is primarily determined by the number of LLM inferences performed. As noted in Section 4.1, P-SHAP and CaptumLIME were allocated the same or slightly more inferences than MExGen in our experiments. Even so, MExGen consistently delivered more faithful explanations.
4.4 Comparison with LLM Self-Explanation
Dataset Model Scalarizer MExGen C-LIME MExGen L-SHAP MExGen LOO Self
XSUM Granite-3.3 Prob 18.9 19.0 18.9 9.5
BART 15.4 14.4 14.2 12.4
DeepSeek-V3 BART 12.7 12.3 12.3 10.5
CNN/DM Granite-3.3 Prob 17.3 17.4 16.9 7.1
BART 11.9 11.0 11.0 8.9
DeepSeek-V3 BART 14.1 14.0 13.5 13.5
Table 2: Areas under the perturbation curve (AUPC) up to $20$% of tokens for the comparison with LLM self-explanation. The scalarizer in the “Scalarizer” column is used for both explanation and evaluation.
LLMs are capable of explaining their own outputs, including by providing numerical attributions to their inputs (Huang et al., 2023b; Madsen et al., 2024). In this experiment, we compare the fidelity of these self-explanations to that of MExGen. Specifically, since the perturbation curves that we use to evaluate fidelity depend only on the ranking of input units, we prompt the LLM to rank units in order of importance to the output that it generated.
LLMs and Datasets
As a test of current LLMs’ ability to self-explain, we chose a large and powerful open-weights LM, DeepSeek-V3. At the time of writing, we could not reliably obtain log probabilities from DeepSeek-V3, so we were unable to apply MExGen with the Log Prob scalarizer and could only use a text-only scalarizer. For this reason, we tested a second LLM, Granite-3.3-8B-Instruct, for which both Log Prob and text-only scalarizers were possible. We used the summarization datasets XSUM and CNN/DM.
Explanation Methods
For self-explanation, we describe in Appendix C.4 the prompt that we used and other details. The most important point is that we made the ranking task easier by following Zhang et al. (2024) in numbering input units with tags and only asking for a list of tags. MExGen was run in the same way as before (see Appendix C.3).
Results
Table 2 shows the AUPC values for this self-explanation experiment. The corresponding perturbation curves are in Figure 12. In all cases, MExGen is more faithful to the LLM’s behavior (as measured by higher AUPC) than the LLM’s self-explanations. Considering Granite-3.3, the advantage of MExGen is especially large (AUPC is double or more) when it uses the Log Prob scalarizer, and less so with the text-only BART scalarizer. The larger DeepSeek-V3 model narrows the gap further on the CNN/DM dataset. Overall, our results indicate that while LLM self-explanations can be good, they are outperformed by algorithms that systematically quantify importance. It is also important to note that this experiment only assesses ranking ability, whereas MExGen also provides real-valued scores, a task that would be harder for an LLM.